What is the pH of a solution containing 0.200 mol sodium acetate and 0.100 mol acetic acid dissolved in 100.0 mL of water, given that the pKa of acetic acid is 4.75?

This amino acid has an overall net charge of -1 at pH 8.5
Select one:
a. arginine
b. cysteine
c. histidine
d. lysine
e. proline
1
Expert's answer
2020-09-24T08:14:30-0400

1

PH = PKa + log ( concentration of sodium acetate / concentration of acetic acid ) .


PH = 4.75 + log 0.2 /100 / 0.1 /100 = 4.75 + log2 = 4.75 + 0.301 = 5.051 Answer .


2.

Answer is" histidine " .